suppose that the demand equation for delmonte, amonopolist is given by p=400-20Q cost equation is given by TC=500+20Q2.find the profit maximizing price and output for the monopolistic
1
Expert's answer
2019-03-20T12:48:22-0400


The profit maximization condition is equality MC = MR.

Find the MS and MR from these equations;

TR=Q*P=400Q-20Q2

MR=(TR)'=400-40Q

MC=(TC)'=40Q

40Q=400-40Q

80Q=400

Q=5

P=400-20*5=300